OpenStudy (anonymous):

Can someone show me the steps to solve this problem? Point A (5,7) is 10 units away from point be (13,y). Solve for the value of y.

OpenStudy (anonymous):

\[\sqrt{\left( 13-5 \right)^{2}+\left( y-7 \right)^{2}}=10\] \[64+\left( y-7 \right)^{2}=100\]

OpenStudy (anonymous):

\[\left( y-7 \right)^{2}=100-64=36\]

OpenStudy (anonymous):

\[y-7=\pm 6\]

OpenStudy (anonymous):

solve for y and get two points.

OpenStudy (anonymous):